OpenStudy (anonymous):

Find the distance between the points: (2,0) and (3,-3)

OpenStudy (anonymous):

use the distance formula: \(\large Distance=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2} \)

OpenStudy (anonymous):

did you try it yet?

OpenStudy (anonymous):

totally lost brother

OpenStudy (anonymous):

oh is it 10

OpenStudy (anonymous):

ok... subtract the x-coordinates... then square that answer... what u got?

OpenStudy (anonymous):

what is the particular solution of ydx + (2x +3y)dy=o when the answer is xy^2 + y^3

OpenStudy (anonymous):

dpalnc i got 10

OpenStudy (anonymous):

is that your answer to my last question or is 10 the distance?

OpenStudy (anonymous):

10 is the distance

OpenStudy (anonymous):

that is homogeneous differential equation

OpenStudy (anonymous):

i think 10 is just the ths part right? : \(\large (x_2-x_1)^2+(y_2-y_1)^2 \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so put a square root over that 10 and that's your answer...

OpenStudy (anonymous):

3.16

OpenStudy (anonymous):

about 3.16..... but exactly \(\large \sqrt{10} \)